OpenStudy (anonymous):

what is the approximate value of log 25? 6 a) .233 b) .557 c) .620 d) 1.796

OpenStudy (anonymous):

any one?

OpenStudy (anonymous):

what is the expression?

OpenStudy (schrodingers_cat):

Is it base 6?

OpenStudy (anonymous):

it it \[\log _6 25\text{ or }\log \frac{25}{6}\,?\]

OpenStudy (isaiah.feynman):